how do you simplify 42xy/49x?
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

42/49 = (6*7)/(7*7) = 6/7

(xy)/x = y


So 42xy/49x = (6y)/x
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------

Answer by stanbon(75887)

how do you simplify 42xy/49x?
---
Ans: cancel common factors:
---
= (6*7*xy)/(7*7x)
-----
= (6y)/7)
-----
= (6/7)y
==================
Cheers,
Stan H.
===========